Status Report

NASA ISS On-Orbit Status 14 June 2008

By SpaceRef Editor
June 14, 2008
Filed under , , ,
NASA ISS On-Orbit Status 14 June 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – off-duty day for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.

STS-124/Discovery returned to Earth this morning after 13d 18h 13min in space, touching down at KSC on the first landing opportunity at 11:15am EDT, after 217 orbits & 5.7 million miles. [During the ISS 1J mission, executed with stellar perfection, its seven-member crew conducted three EVAs, delivered & installed the JEM (Japanese Experiment Module) Kibo with its RMS (Robotic Maneuvering System), brought up new Expedition 17 crewmember Gregory Chamitoff and returned his predecessor Garrett Reisman who spent 95d 8h 47m in space (with ~89d on board ISS). It was the 123rd flight of a Space Shuttle, the 26th Shuttle mission to visit the station, the 35th for Discovery and the 69th landing at KSC. Welcome back, Discovery! Next up: STS-125/Atlantis on 10/8, on Service Mission 4 (SM4) to the Hubble Space Telescope (HST).]

The FE-1 serviced the Russian BMP (Harmful Impurities Removal System), starting the "bake-out" cycle to vacuum on absorbent bed #2 of the regenerable dual-channel filtration system. The regen process will be terminated tonight at ~5:15pm EDT. Filter bed #1 was regenerated yesterday. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle is regularly done every 20 days.]

Having reached Day 15 of his flight, Greg Chamitoff began his first session with the NASA/JSC experiment NUTRITION w/Repository, for which he had to forego exercising and food intake for eight hours. Today’s protocol consisted of blood draws. Later tonight, the FE-2 will also set up the equipment for the 24-hour urine collections which start with the first void early tomorrow morning and continue through Monday morning. [After performing self-phlebotomy, i.e., drawing blood samples (from an arm vein), the samples were first allowed to coagulate in the Repository for 20-30 minutes, then spun in the HRF RC (Human Research Facility/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). No thruster activity was allowed during the blood drawing. The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned. Background: NUTRITION is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight; this includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes. The Clinical Nutritional Assessment profile currently required on all U.S. Astronauts collects blood and urine samples preflight and postflight. NUTRITION expands this protocol by also capturing inflight samples and an additional postflight sample. Furthermore, additional measurements are included for samples from all sessions, including additional markers of bone metabolism, vitamin status, and hormone and oxidative stressor tests. The results will be used to better understand the impact of countermeasures (exercise and pharmaceuticals) on nutritional status and nutrient requirements. The Clinical Nutritional Assessment profile (MR016L), first started on two Mir crewmembers and then on all ISS US crews, nominally consists of two pre-flight and one post-flight analysis of nutritional status, as well as an in-flight assessment of dietary intake using the FFQ (Food Frequency Questionnaire). The current NUTRITION project expands MR016L testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.]

The crew performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

Volkov & Kononenko also performed regular maintenance inspection & cleaning on fan grilles in the FGB (TsV & TsV2), SM (VPkhO & VPrK), DC1 (V3), and COL (Columbus Orbital Laboratory) for ESA.

Additionally, CDR Volkov conducted the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables. [Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.]

Greg Chamitoff filled out the regular FFQ (Food Frequency Questionnaire), his first, on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

At ~9:15am EDT, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1).

Afterwards, Greg transferred the crew’s exercise data file to the MEC for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

The Russian crewmembers had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Oleg at ~7:00am, Sergey at ~8:40am.

For CDR Volkov, the Russian discretionary task list held one job for today – continuing the GFI-8 "Uragan" (hurricane) earth-imaging session begun yesterday, using the D2X with the F800 tele-lens. [Uplinked target areas were the coastal line of Lesser Aral Sea, the northern section of the sea separated by a dam from the Greater Aral, Katunsky preserve inside of Katun river bend enveloping the mountains, South shores of Teletsk Lake, woodland area of Sayan mountains, Irkut river valley running in direction of Baikal, north shores of Lake Baikal near Angara, Trans-Baikal woodland area, woodlands near Russian population areas along Amur and Ussuri rivers, ocean shores, upper reaches of Mzymta river valley from Adler, Caucasus mountain range near Teberda-Donbai, flooded Volga-Akhtubinsk alluvial plain after water reservoir releases, oil spills on the north of Caspian Sea, the Yasnyi Cosmodrome left of track, Katun river valley, the western shore of Teletsk Lake in nadir woodland area of Sayan mountains, general view of Carpathian mountains, the valley of one of the rivers in Moldova, the Don river valley, wooded areas in Orenburg steppe, general view of the Ural mountains, the Irtysh river valley, general views of the Alps, general views of Byelorussia, Ugra National Park, Don river valley, general views of the Aral Sea, and the Kyrgyz mountain range to the west of Issyk-Kul.]

Also working from the “time permitting” task list, FE-1 Kononenko conducted another run, his sixth, of the Russian DZZ-2 "Diatomeya" ocean observations program, using the NIKON-F5 DCS and the HDV (high-definition) video camcorder from SM windows 8 for ~20 min to record high production water areas for obtaining data on color field composition in dynamic regions of the ocean and in near-estuary areas of large rivers, plus the current cloud cover above these waters, its rating, and special forms of cloud formation. [Target zones today were in the Pacific Ocean (Tokelau Islands – Californian upwelling).]

A second discretionary task list job for Oleg was a session for Russia’s Environmental Safety Agency (EKON), making observations and taking KPT-3 aerial photography of environmental conditions of Astrakhan using the Nikon D2X digital camera with the SIGMA 300-800mm telephoto lens.

ITCS Temporary Leak Event: Yesterday a small leak occurred in a QD (quick disconnect) when the crew disconnected the ITCS LTL (Internal Thermal Control System/Low Temperature Loop) jumper to the AR (Air Revitalization) in support of the standard CDRA (Carbon Dioxide Removal Assembly) deactivation procedure. A second, larger leak occurred when the disconnect was repeated on ground advise. The AR Rack return side QD was subsequently disconnected with no leak occurring. The total accumulator quantity drop (all in the LTL) was about 18%, approximately 2 liters. The ITCS coolant is Toxicity-0, and the crew cleaned up completely after the leak, discarding the towels in a wet trash bag (if the coolant liquid is allowed to evaporate, the remaining dry biocide residue becomes Toxicity-1, possibly a mild irritant).

Weekly Science Update (Expedition Seventeen — Week 8)

PADLES (Passive Dosimeter for Lifescience Experiment in Space): Twelve passive dosimeters were set up into JEM at different positions (6/12). [Area PADLES measures radiation exposure levels onboard the ISS using passive & integrating dosimeters to detect radiation levels. These dosimeters are located near the biological experiment facilities and on the end of Kibo.]

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): In progress.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. Its power feeder #1 has been de-activated to support the JEM install activities (EVA-2). After the EVA-2, the EuTEF instruments have then been re-activated nominally. A DHPU (Data Handling Processing Unit) software patch is currently developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments.- DEBIE-2: Link error still under investigation. Some short-duration science measurements are on going since 6/11;- DOSTEL: On-going science acquisition.- EuTEMP: Currently inactive as planned.- EVC: Right after the EuTEF platform re-activation, EVC was successfully powered on and acquired very good images of Europe. Unfortunately, further new attempts were not successful early this week.- EXPOSE: On-going science acquisition.- FIPEX: A new script has been started after the re-activation of the EuTEF platform. The instruments sensors went off on 6/6, and the measurement sequence had to be restarted, no significant science impact. Now acquiring science. – MEDET: Acquiring science since 6/5. – PLEGPAY: Several measurements have been performed with the Langmuir probe instrument during the STS-124 (1J) docking maneuver. The instrument is now powered on, but not in science acquisition mode. Unfortunately, no Langmuir probe measurements could be planned in conjunction with the Shuttle undocking.- TRIBOLAB: The instrument has been put in Thermal Stabilization Mode after EuTEF power feeder #1 re-activation. Pin on Disk Run #4 (POD#4) experiment was successfully started on 6/12.

FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. Further troubleshooting activities with CEM-U(pper) Optical Module are required.

GEOFLOW: A new lamp set is delivered by STS-124 (1J). The start of GEOFLOW is pending further FSL troubleshooting and remaining commissioning activities (check of FSL optical modes).

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: “Greg, thanks for all your work and dedication in obtaining your early Integrated Immune samples. They will provide us with valuable, early in-flight data that will help us asses the risk of immune deregulation in spaceflight.”

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): In progress.
SOLAR (Solar Monitoring Observatory): A new Sun observation period has started on 6/3. After the JEM activation, SOLAR power feeder #1 has been reactivated, and we are now acquiring science nominally. Some criss-cross maneuvers have been successfully performed on 6/8 & 6/12. These activities aim at correlating the sun sensor positions on the SOLAR Coarse Pointing Device (CPD) and the SOVIM / SOLSPEC instruments. The SOLACES instrument was also used during these activities. The present Sun observation window will close on 6/16. – SOVIM: Acquiring science. – SOLSPEC: Acquiring science.- SOLACES: Acquiring science.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

3-D SPACE : PI (Principal Investigator)/crew conference was completed on 6/12. First session was performed (nominally, but with more crew time than planned) on 6/13. “Thank you for your efforts in getting the data for 3D-Space!”

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The four ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BLB (BIOLAB) could not be activated from ground yet, pending resolution of the Smoke Detector issue. After the 1J Flight, the remaining WAICO-#1 ECs will be disposed and replaced by six Reference ECs on Rotor B.

CEO (Crew Earth Observations): Ongoing.

No CEO (Crew Earth Observation) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

ISS Orbit (as of this morning, 6:18am EDT [= epoch]):
Mean altitude — 339.2 km
Apogee height — 343.0 km
Perigee height — 335.4 km
Period — 91.32 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005682
Solar Beta Angle — -18.4 deg (magnitude increasing)
Orbits per 24-hr. day — 15.77
Mean altitude loss in the last 24 hours — 35 m
Revolutions since FGB/Zarya launch (Nov. 98) — 54809

Significant Events Ahead (all dates Eastern Time, some changes possible.):
06/19/08 — ATV1 Reboost
07/10/08 — Russian EVA-20 (7/10-11)
09/05/08 — ATV1 undocking
09/09/08 — Progress M-64/29P undocking (from FGB nadir)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (SM aft port)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir)
11/03/08 — Soyuz TMA-13/17S relocation
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
12/04/08 — STS-119/Discovery/15A launch – S6 truss segment
12/06/08 — STS-119/Discovery/15A docking
12/15/08 — STS-119/Discovery/15A undocking
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/??/09 — Six-person crew on ISS (following Soyuz 18S-2 docking)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.